OpenStudy (anonymous):

Can't find part B! For the sequence An = 8/3^n, A) its 5th partial sum S_5=968/243 B)its nth partial sum Sn = ?

jimthompson5910 (jim_thompson5910):

I'm assuming the sequence is \[\large A_{n} = \frac{8}{3^n}\]

OpenStudy (anonymous):

That is correct.

jimthompson5910 (jim_thompson5910):

Break it up and rewrite it to get \[\large A_{n} = \frac{8*1}{3^n}\] \[\large A_{n} = \frac{8*1}{1*3^n}\] \[\large A_{n} = \frac{8}{1}*\frac{1}{3^n}\] \[\large A_{n} = 8*\frac{1}{3^n}\] \[\large A_{n} = 8*\frac{1^n}{3^n}\] \[\large A_{n} = 8*\left(\frac{1}{3}\right)^n\]

jimthompson5910 (jim_thompson5910):

What do you notice about the last equation?

OpenStudy (anonymous):

I'm not sure, it just seems reminiscent of a the formula used to find Sn...or one of the others.

jimthompson5910 (jim_thompson5910):

actually, let me rewrite it in a slightly different way one sec

jimthompson5910 (jim_thompson5910):

\[\large A_{n} = \frac{8*1}{3^n}\] \[\large A_{n} = \frac{8*1}{1*3^n}\] \[\large A_{n} = \frac{8}{1}*\frac{1}{3^n}\] \[\large A_{n} = 8*\frac{1}{3^n}\] \[\large A_{n} = 8*\frac{1^n}{3^n}\] \[\large A_{n} = 8*\left(\frac{1}{3}\right)^n\] \[\large A_{n} = 8*\left(\frac{1}{3}\right)\left(\frac{1}{3}\right)^{n-1}\] \[\large A_{n} = \frac{8}{3}\left(\frac{1}{3}\right)^{n-1}\]

jimthompson5910 (jim_thompson5910):

What form is that last equation in?

OpenStudy (anonymous):

I'm sorry, I'm really missing this...

jimthompson5910 (jim_thompson5910):

what sequences have you studied so far?

OpenStudy (anonymous):

My teacher has just started on arithmetic series and sequences and he mentioned geometric ones.

jimthompson5910 (jim_thompson5910):

what form is an arithmetic sequence in

OpenStudy (anonymous):

\[a_n=a_1+(n-1)d?\]

jimthompson5910 (jim_thompson5910):

good, how about geometric

OpenStudy (anonymous):

\[a_k+1=a_kr\]

OpenStudy (anonymous):

With the one subbed a long with the k.

jimthompson5910 (jim_thompson5910):

do you know of any other way to express a geometric sequence in general?

OpenStudy (anonymous):

I'm not sure if it pertains, but I had read in my book of : xn = ar(n-1)

OpenStudy (anonymous):

r^(n-1)*

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

look at the last equation I wrote above notice how it matches up with that form you just wrote

OpenStudy (anonymous):

Yes, it certainly does.

jimthompson5910 (jim_thompson5910):

so what are 'a' and 'r'

OpenStudy (anonymous):

a = 8/3 and r = (1/3)^(n-1)

jimthompson5910 (jim_thompson5910):

close, a = 8/3 is correct

jimthompson5910 (jim_thompson5910):

but r = (1/3)^(n-1) is NOT correct

OpenStudy (anonymous):

Is it because r stands for common ratio? Because if so, it'd be (1/2)^(n-1)

jimthompson5910 (jim_thompson5910):

it should be r = 1/3 do you see how/why?

OpenStudy (anonymous):

Yeah, I think I understand why.

jimthompson5910 (jim_thompson5910):

now that you know 'a' and 'r', how can you use them to find the nth partial sum?

OpenStudy (anonymous):

Is there actually a formula for that?

jimthompson5910 (jim_thompson5910):

yes, one sec

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

check out this page http://www.regentsprep.org/regents/math/algtrig/ATP2/GeoSeq.htm

OpenStudy (anonymous):

I see the Sn formula, if I were to plug in my values, it should look something like this:

OpenStudy (anonymous):

\[(8/3 (1-1/3)^n)/(1-1/3)\]

OpenStudy (anonymous):

Which is 2^(n+2)3^-n

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

ok this is what I'm getting \[\large S_{n} = \frac{a(1-r^n)}{1-r}\] \[\large S_{n} = a\frac{1-r^n}{1-r}\] \[\large S_{n} = \frac{8}{3}\frac{1-(1/3)^n}{1-1/3}\] \[\large S_{n} = \frac{8}{3}\frac{1-(1/3)^n}{2/3}\] \[\large S_{n} = \frac{8}{3}(1-(1/3)^n)*\frac{3}{2}\] \[\large S_{n} = \frac{8}{3}*\frac{3}{2}(1-(1/3)^n)\] \[\large S_{n} = \frac{8*3}{3*2}(1-(1/3)^n)\] \[\large S_{n} = \frac{24}{6}(1-(1/3)^n)\] \[\large S_{n} = 4(1-(1/3)^n)\]

OpenStudy (anonymous):

Why did you move the a out to the front though?